OpenStudy (anonymous):

A lamp hangs from the ceiling at a height of 2.9 m. If the lamp breaks and falls to the floor, what is it's impact speed?

OpenStudy (anonymous):

A)61.2m/s B)56.8m/s C) 7.5m/s D)14.2m/s

OpenStudy (shamim):

U know the formula V^2=u^2+2*g*h Given u=0 h=2.9m

OpenStudy (shamim):

Tell me v=?

OpenStudy (anonymous):

Its asking for the impact speed. Doesnt give the velocity @shamim

OpenStudy (shamim):

Wt do u mean by impact speed?

OpenStudy (anonymous):

@perl HELP lol

OpenStudy (perl):

@janello shamim is right

OpenStudy (perl):

you can use the formula v^2=u^2+2*g*h v = final speed (here speed of impact) u = initial speed (here zero) g = acceleration due to gravity (9.8 m/s^2) h = height (here 2.9 meters)

OpenStudy (anonymous):

The answer i get isnt one of the amswer choices.

OpenStudy (anonymous):

THANK YOU SO MUCH @perl